What is the probability of rolling a 6-sided dice and getting an odd number three times and a row?
The cube has 6 faces, so it can come up in 6 different ways.
Three of the possible outcomes are odd numbers. (1, 3, and 5)
Probability of an odd number on the first roll = 3/6 = 1/2 .
Probability of an odd number on the second roll = 3/6 = 1/2 .
Probability of an odd number on the third roll = 3/6 = 1/2 .
Probability of an odd number on all 3 rolls = (1/2) x (1/2) x (1/2) = 1/8 = 0.125 = 12.5%
In a multiple choice question the correct answer will be found?
... most easily if you have some knowledge of the subject,
for example by having studied it.
First white = 10/20
second white = 9/19
third white = 8/18
fourth white = 7/17
firth white = 6/16
Total probability = 10/20 x 9/19 x 8/18 x 7/17 x 6/16
= 21/1292
~= 0.0163
2, 4, and 6 are multiples of 2. This is 1/2 of the face, so the probability is 1/2
The number of dots on a regular die?
There are six faces numbered 1 to 6 so:
6 + 5 + 4 + 3 + 2 + 1 = 21
What is a belief that an event will occur?
something that some people believe that in 2012, the world will end.
is that what your trying to ask?
If you roll a nine-sided die what is the probability you will roll a value greater than zero?
If zero is the number on one of the 9 sides, then the probability is 8/9 = 88.9% (rounded)
If zero is not the number on any one of the 9 sides, then the probability is 100%.
Each cube can land in 6 different ways.
Two cubes can land in (6 x 6) = 36 different ways.
There are 6 ways that succeed:
3 - 4
4 - 3
3 - 5
5 - 3
3 - 6
6 - 3
The probability is (6/36) = 1/6 = 16 and 2/3 percent
The probability it's a king is 1/13.
The probability it's a heart is 1/4.
The probability that it is a king OR a heart is 4/13, or about 31%
There are 4 kings, and 13 hearts, but the king of hearts is both, so you can't count that twice, so you have to subtract one.
There is the king of hearts; 3 kings that are not hearts, and 12 hearts that are not the king. 1 + 3 + 12 = 16.
Divide that by the total number of cards in the deck (52), = 16/52, which reduces to 4/13. In percentage, the probability is about 31%.
What are the odds of getting mauled by a bear?
You stand a better chance of being struck by lighting than mauled by a bear. According the Ontario Ministry of Natural Resources there are about 10 to 20 bear attacks in North America every year, with about 2 or 3 deaths.
Not sure what the odds are, but given the huge number of people that go tromping through the forest and the close proximity man and bear have to one another, the odds are pretty low.
What numbers in the lottery have the most probability of winning?
They all have the same probability, even though in the past some have been drawn more often than others. It's still completely random though!
Multiple choice questions and answers on DNS?
You can post your individual math problems on WikiAnswers, but you cannot ask for answer keys for specific educational programs.
See related link.
One in four.
The possibilities are;
Red Red
Blue Red
Red Blue
Blue Blue
That means the number of ways an event can occur is one, while the number of possible outcomes is four. The definition of probability is the number of ways the event can occur over the number of possible outcomes, hence 1/4
What is the probability of tossing a coin 25 times and it landing on heads twice?
75 over 2^23 (2 times itself 23 times, 2x2x2x2x2 etc...)
Normal curve is the meaning of standard normal distribution?
No, the normal curve is not the meaning of the Normal distribution: it is one way of representing it.
In a multiple choice test what is the probability of getting 4 correct answers?
That's really going to be heavily dependent on whether or not you know something about
the subject material, and also on the total number of questions on the test.
It has even been known to happen from time to time that there are people who will score
100% on one of those, i.e. all correct answers.
Since that happens at a rate that far exceeds random probability, many educators and
administrators have begun to suspect that some of these test-takers may actually be
going so far as to study the material in advance! But the allegation hasn't been proven.
King of Hearts (1/52)x Queen of Hearts (1/51)=1/2652
How do you calculate a continuous joint probability distribution?
You simply have a function with two (or more) arguments which are continuous.
For example, z = p(x,y) would be a surface in 3 dimensions where x and y are the values taken by the two variables, X and Y respectively, and z is the probability associated with them.
w = g(x,y,z) would be a hyper-surface in 4-dimensional space and so on.
What amount of sampling error is acceptable in a national sample of 1500 to 3000 respondents?
There can be no set value.
An acceptable level of sampling error for a company making high precision machine parts is likely to be very different from the sampling error for household incomes, for example.
If the paired differences are normal in a test of mean differences, then the distribution used for testing is the
What is the probability of tossing a 2 or 3 on a number cube?
Each face has probability 1/6, since it's one face out of the possible 6.
So the probability of either of two faces is 2/6, or 1/3.
If you roll a die what is the probability that a one will appear on top?
Depends, but if you're talking theoretically the answer would be 1:6